# Privacy Perspectives and Practices of Chinese Smart Home Product Teams

### v1

@@ Abstract
Previous research has explored the privacy needs and concerns of device owners, primary users, and different bystander groups with regard to smart home devices like security cameras, smart speakers, and hubs, but little is known about the privacy views and practices of smart home product teams, particularly those in non-Western contexts. This paper presents findings from 27 semi-structured interviews with Chinese smart home product team members, including product/project managers, software/hardware engineers, user experience (UX) designers, legal/privacy experts, and marketers/operation specialists. We examine their privacy perspectives, practices, and risk mitigation strategies. Our results show that participants emphasized compliance with Chinese data privacy laws, which typically prioritized national security over individual privacy rights. China-specific cultural, social, and legal factors also influenced participants’ ethical considerations and attitudes toward balancing user privacy and security with convenience. Drawing on our findings, we propose a set of recommendations for smart home product teams, along with socio-technical and legal interventions to address smart home privacy issues—especially those belonging to at-risk groups—in Chinese multi-user smart homes.
@@ 1 Introduction
Smart home devices have seen widespread global adoption, with projections estimating approximately 311.9M smart homes worldwide by 2027 [69]. This rapid adoption has significantly raised privacy concerns among different stakeholders, including device owners, primary users, secondary users, and bystanders in multi-user smart homes (e.g.  [135, 146, 47, 139, 81, 59, 96, 71] [11, 36, 147, 91, 83, 27, 38, 103, 130, 93, 20, 107]). Despite this extensive literature from an end-user perspective, few studies have considered smart home product teams. This is a notable gap, as the perceptions, perspectives, and practices of those designing and developing smart home devices may play a crucial role in multi-user smart home privacy. Although some work has been conducted with smart home designers [24, 25, 12, 13], there remains a gap in understanding how smart home product teams—comprising software/hardware engineers, product managers, legal/compliance advisors, and marketing staff—address privacy-related challenges.
Furthermore, the current body of research on smart home privacy is geographically focused on WEIRD (Western, Educated, Industrialized, Rich, Democratic) countries [89, 24, 25, 102]. Smart home privacy in non-WEIRD countries has received limited scholarly attention [49]. As a non-WEIRD nation, China offers a distinct set of social, cultural, and legal characteristics that can surface novel privacy challenges and design considerations in the realm of smart home technologies [51], including: i) the country’s accelerated digital transformation [31]; ii) the high rate of smart home adoption, with over 78M Chinese households using such devices in 2023 [109]; iii) the large number of smart home product teams in China [106]; iv) the introduction of new privacy-focused regulations within the country (see §2.4); and v) China’s tech landscape—driven by firms like Xiaomi and Huawei instead of Western giants—which has given rise to alternative data governance practices and innovation pathways [143].
Our study is the first to examine the privacy views and practices of Chinese smart home product team members, focusing on those involved in designing, developing, or evaluating smart home devices. These teams typically include software and hardware engineers, UX researchers/designers, product and project managers, legal/privacy experts, and marketing or operations specialists.
Our research questions (RQs) are as follows:
@@
RQ1. Participants highlighted the importance of legal compliance in smart home device design and development, often engaging in routine reviews and showing a strong ethical commitment to user privacy. However, many—particularly UX designers—lacked adequate knowledge of privacy laws, data practices, and formal training. They also noted that China’s unique sociocultural context (e.g., collectivism and Confucian values) contributes to lower privacy concern among Chinese smart home device users and limited legal literacy among product teams, shaping the design of privacy features.
@@
RQ2. Participants encountered several challenges, including: (1) managing complex and dynamic relationships between government authorities, companies, and users - for example, situations where law enforcement agencies could access user data, bypassing internal privacy safeguards within companies; (2) limited attention to the privacy needs of at-risk groups (e.g., older adults, domestic workers), stemming from difficulties in conducting empirical research with these populations and ethical limitations; and (3) the need to balance users’ privacy with UX, such as ease of use and convenience, particularly within the shared environment of multi-user smart homes.
@@
RQ3. To address the challenges identified above, participants used various privacy practices and risk mitigation strategies. These included technical measures such as restricting data access and implementing decentralized data processing. They also advocated for social initiatives to improve user privacy awareness (e.g., community-level outreach activities) and called for clearer legal guidelines in smart home design and development. Although participants acknowledged recent improvements in local privacy legislation, they highlighted a gap between legal frameworks and actual development practices, often caused by the fast-paced evolution of smart home technologies in China.
@@
Contributions. We conducted the first exploratory interview study focusing on Chinese smart home product team members to investigate their perspectives and practices related to privacy. Unlike prior studies, our research involves comprehensive roles within smart home product teams—including product managers, designers, engineers, marketers, and compliance officers—from predominantly larger companies, thereby providing a comprehensive overview (see Table I). Additionally, this study is the first to explore how these product team members navigate and implement privacy laws and ensure legal compliance during smart home product design and development. Our findings provide valuable and rich empirical insights into privacy practices in the Chinese smart home industry, offering implications for product teams, researchers, and policy makers, especially in light of China’s distinct sociocultural environment. Based on our empirical evidence, we extend the discussion on Privacy by Design by emphasizing the necessity of responsible and transparent guidelines adapted to China’s unique sociocultural context, such as Confucian values, collectivist norms, and China’s culturally specific understanding of “privacy” (see §5.3). Further, our findings highlight disparities and power dynamics between larger companies—which typically have greater resources, dedicated compliance teams, and structured internal legal support—and small- and medium-sized enterprises (SMEs), which often face significant challenges in privacy compliance due to resource constraints and limited legal expertise. Based on our empirical evidence, we strongly recommend clearer legal frameworks to guide privacy practices, specifically designed to support compliance efforts in SMEs. Finally, we call for enhanced awareness of privacy issues among both the general public and product team members through educational initiatives in communities and schools, as well as targeted training for legal/regulatory compliance and data management (see §5.4).
@@ 2.1 Multi-User Smart Home Privacy
Previous studies have explored the privacy, security, and safety needs, preferences, and concerns of smart home device owners and primary users [146, 47, 1, 2, 135, 122, 139, 17, 112, 48, 70, 59, 81], including Chinese users [74, 58, 50], emphasizing the importance of data ownership by users, data practice transparency, and informed user consent. The studies have also highlighted how social norms, including privacy norms, along with factors such as power dynamics, social structures, demographics, and technical expertise, influence views on smart home device adoption and use [2, 16, 112, 52, 107]. Other papers have explored the privacy implications of smart home devices for different bystander groups who share a living space, including: visiting, live-in, and uninvolved [99], highlighting privacy and security concerns of at-risk groups such as domestic workers [19, 20, 107, 9, 10, 62, 61, 36, 4, 51] and household members living in hostile environments [15, 72]. Additionally, some studies engaged device owners/users and bystanders in the co-design and evaluation of prototypes for select design interventions proposed in the academic literature [81, 45, 132, 122, 82, 70, 8, 94, 136].
Building on the findings of the aforementioned studies involving primarily device owners and bystanders, researchers proposed technical solutions to improve multi-user smart home privacy, including minimizing data collection for essential functions [10, 81, 84]; selectively limiting data collection in sensitive situations [30, 85, 45, 113]; designing privacy nudges and notifications [41, 70, 139, 8, 140]; implementing temporary data storage mechanisms [9, 84, 45, 30]; using smart home dashboards for data visualization [81, 132, 122, 82, 140]; creating separate user profiles (e.g., guest profiles with limited data access and control) [15, 41, 59, 70, 81, 139]; employing tangible practices (e.g., camera lens lids) [8, 7, 34, 93, 94, 140, 132]; and implementing communication and negotiation mechanisms to state privacy preferences [45, 70, 136, 30, 11, 147]. Other studies proposed social interventions to address privacy concerns in multi-user smart homes, e.g., design guidelines aimed at supporting smart home product teams [19, 24, 11].
@@ 2.2 Privacy in Technology Design/Development
Previous research on privacy in digital technology design and development focused primarily on the challenges encountered by software developers of mobile apps in eliciting and incorporating privacy-related requirements. These challenges include privacy- and security-related issues in development [75, 118, 64, 101, 5], the failure to comply with privacy laws and regulations [14, 117], and guidance documentation that inadvertently leads developers to privacy-compromising decisions [119, 66]. In addition, privacy concerns have been identified in APIs and tools used by programmers [29, 44]. Research also showed that developers often deprioritize privacy due to stakeholder pressures [37, 66], placing greater emphasis on revenue generation and functionality than on privacy protection  [120, 65, 114]. (Refer to [89] for a review of the factors influencing software developers’ security and privacy practices.) Unlike prior studies, our paper investigates the privacy perceptions and practices of Chinese smart home product teams, hence extending the focus from software developers to a broader range of product team members across diverse roles and organizations, focusing on smart home devices.
@@
In smart homes, Chalhoub et al. examined how product teams at UK companies integrate UX principles into the security & privacy design of smart home cameras, proposing solutions such as customizable privacy settings and transparent user consent mechanisms [24, 25]. These studies were conducted in a WEIRD context, focused exclusively on one smart home device: smart cameras, and focused primarily on end-user experience. They did not study how product teams perceived the privacy concerns of other stakeholders in smart homes, including bystanders, and did not take into account the power dynamics that exist among these groups. Zhang and Xie have examined the main activities of and challenges faced by IoT developers—mainly hardware engineers, embedded software developers, and app developers—in Chinese (non-WEIRD) start-ups [145]. Although the study did not focus on security and privacy in the smart home context, and recruited participants only from startups, it found that participants frequently overlooked security issues due to limited resources and skills.
@@
Finally, recent work [10, 11] showed how socioeconomic factors, social norms, and religious beliefs influence power dynamics in Jordanian (non-WEIRD) smart homes. However, these studies did not explore the privacy perceptions and practices of local smart home product teams, focusing on end-users and a bystander group: domestic workers. While follow-up work in Jordan [12] has focused on smart home designers and their lack of privacy-support tools and guidelines, our study offers novel insights into Chinese product teams’ perspectives and practices, examining privacy considerations across multiple roles. Additionally, our research highlights how cultural values (e.g., filial piety), priorities (e.g., prioritizing legal compliance over individual privacy), and hierarchical, collectivist dynamics influence privacy trade-offs within these teams.
@@ 2.3 Confucianism and Privacy in China
Confucianism, a foundational philosophy in Chinese culture, emphasizes social harmony, hierarchical relationships, and the importance of family and community [134]. Central to Confucian thought are concepts like filial piety (“xiao”), which underscores respect and duty toward one’s parents and elderly people, and collectivism, which prioritizes group needs over individual desires [54]. Some studies have highlighted that traditional Confucian thinking, which significantly affects the understanding of privacy within the Chinese family unit, is often associated with the indigenous Chinese concept of “yin si” with notions of unspeakable immorality, indecency, or shame, providing minimal support for the development of legal norms that favor individual privacy protection [87, 128, 138]. Consequently, privacy protection norms in China have developed at a slower pace when compared to Western societies, where privacy is deeply rooted in concepts of liberal individualism [18] and central to liberal political theory [53].
2.4 Legal/Regulatory Frameworks in China
@@
In China, companies’ privacy practices are primarily governed by the Personal Information Protection Law (PIPL) [^1] Personal Information Protection Law (PIPL), the Cybersecurity Law (CSL) [^2] Cybersecurity Law (CSL) of the People’s Republic of China, Data Security Law (DSL) [^3] Data Security Law (DSL) of the People’s Republic of China, and the Provisions on Online Protection of Children’s Personal Information (POPCPI) [^4] Provisions on Online Protection of Children’s Personal Information. Appendix B details the similarities and differences between these privacy laws and regulations. These legal frameworks provide data privacy protection and the scope of state oversight over data processors and companies [63, 126]. CSL and DSL contain provisions related to personal information protection but they both prioritize national security over individual privacy rights [22]. In contrast, PIPL focuses on individuals’ privacy within the Chinese context (e.g., Chapter 3). However, certain terms in the PIPL—such as separate consent and the criteria for lawful anonymization—are open to interpretation, which can create uncertainty in how they are applied [137, 125]. PIPL supports individual data protection, but allows government access under certain circumstances (e.g., Chapter 2, Section 3). Furthermore, enforcing data protection laws in China is complex due to the involvement of multiple administrative authorities, overlapping responsibilities, inconsistent enforcement practices, and internal contradictions within the legal framework [32, 144].
@@
Age Gender Education Experience Location Role Firm size Type Designed & Developed device(s)
@@
P1-EN 25-34 Man Master 6-10 yrs Shanghai Hardware engineer 1000+ MNC Smart health devices/monitors
@@
P2-PM 25-34 Man Master 1-5 yrs Shanghai Product manager 1000+ MNC Smart speakers, Smart home cameras
@@
P3-PM 25-34 Man Bachelor 6-10 yrs Shanghai Product manager 11-50 PE Baby cameras/monitors, Smart lighting
@@
P4-PM 25-34 Man Bachelor 1-5 yrs Guangdong Product manager 1000+ PE Baby cameras/monitors, Smart home cameras, Door locks/doorbells
@@
P5-UX 25-34 Man Bachelor 10+ yrs Shanghai UX design manager 1000+ PE Smart speakers, Smart home cameras, Door locks/doorbells, Smart thermostats, Smart TVs, Smart lighting, Smart plugs
P6-UX 25-34 Woman Master 6-10 yrs Sichuan UX designer 1000+ SOE Smart speakers, Smart TVs
@@
P7-PM 25-34 Man Master 1-5 yrs Guangdong Product manager 501-1000 PE Smart speakers, Door locks/doorbells
@@
P8-UX 25-34 Woman Master 6-10 yrs Shanghai UX designer 1000+ MNC Smart health devices/monitors, Smart thermostats
@@
P9-UX 25-34 Man Master 1-5 yrs Chongqing UX designer 1000+ SOE Smart speakers, Smart home cameras, Smart thermostats
@@
P10-PM 35-44 Man Bachelor 10+ yrs Jiangsu Product manager director 1000+ PE Smart speakers, Smart home cameras, Door locks/doorbells, Smart health devices/monitors, Smart thermostats, Baby cameras/monitors, Smart lighting, Smart plugs, Smoke detectors/alarms
P11-PM 25-34 Man Master 6-10 yrs Zhejiang Project manager 1000+ SOE Smart home cameras
@@
P12-PM 25-34 Woman Master 1-5 yrs Beijing Product manager 1000+ PE Smart home cameras, Door locks/doorbells, Baby cameras/monitors
@@
P13-UX 25-34 Man Master 6-10 yrs Zhejiang UX designer 1000+ SOE Smart home cameras, Door locks/doorbells, Smoke detectors/alarms
@@
P14-EN 25-34 Man Bachelor 6-10 yrs Guangdong Sr. software engineer 51-100 PE Smart speaker, Smart home cameras, Door locks/doorbells, Smart health devices/monitors, Smart Thermostats, Smart lighting, Smart plugs, Smoke detectors/alarms, Kitchen appliances
@@
P15-EN 35-44 Man Master 10+ yrs Shannxi Software director 1000+ SOE Smart health devices/monitors, Smart TVs, Kitchen appliances
@@
P16-PM 25-34 Woman Master 1-5 yrs Guangdong Product manager 0-10 PE Smart speakers, Smart home robots
@@
P17-EN 25-34 Man Doctoral 1-5 yrs Beijing Algorithm engineer 501-1000 PE Smart watch, Smart home cameras
@@
P18-LE 25-34 Man Master 1-5 yrs Beijing Safety compliance engineer 1000+ PE Smart TVs, Smart Thermostats, Smart lighting, Smart speaker, Smart plugs, Smart home cameras, Door locks/doorbells, Smart health devices/monitors, Smoke detectors/alarms
@@
P19-LE 25-34 Woman Master 1-5 yrs Shanghai Compliance officer 1000+ PE Smart home robots
@@
P20-LE 25-34 Man Bachelor 6-10 yrs Guangdong Product compliance engineer 101-500 PE Smart plugs, Smart home cameras, Door locks/doorbells, Kitchen appliances
P21-EN 25-34 Man Master 6-10 yrs Guangdong Software engineer 1000+ SOE Smart home cameras
@@
P22-EN 25-34 Woman Bachelor 1-5 yrs Jiangsu Software engineer 1000+ PE Kitchen appliances, Smart home robots
@@
P23-LE 25-34 Woman Master 1-5 yrs Zhejiang Compliance officer 1000+ SOE Smart speaker, Smart home cameras, Baby cameras/monitors, Kitchen appliances
@@
P24-EN 35-44 Woman Bachelor 10+ yrs Chongqing Software director 1000+ SOE Smart thermostats, Smart speakers, Smart home cameras
@@
P25-MS 25-34 Woman Master 6-10 yrs Shanghai Marketing executive 101-500 MNC Smart plugs, Kitchen appliances
@@
P26-MS 25-34 Woman Master 6-10 yrs Shanghai Digital marketing specialist 51-100 PE Smart home cameras

### v2

@@ Abstract
Previous research has explored the privacy needs and concerns of device owners, primary users, and different bystander groups with regard to smart home devices like security cameras and smart speakers, but little is known about the privacy views and practices of smart home product teams, particularly those in non-Western contexts. This paper presents findings from 27 semi-structured interviews with Chinese smart home product team members. We examine their privacy perspectives, practices, and risk mitigation strategies. Our results show that participants emphasized compliance with Chinese data privacy laws, which typically prioritized national security over individual privacy rights. China-specific cultural, social, and legal factors also influenced participants’ ethical considerations and attitudes toward balancing user privacy and security with convenience. Drawing on our findings, we propose a set of recommendations for smart home product teams, along with socio-technical and legal interventions to address smart home privacy issues—especially those belonging to at-risk groups—in Chinese multi-user smart homes.
@@ 1 Introduction
Smart home devices have seen widespread global adoption, with projections estimating approximately 672.5 million users of smart home devices worldwide by 2027 [62]. This rapid adoption has significantly raised privacy concerns among different stakeholders, including device owners, primary users, secondary users, and bystanders in multi-user smart homes (e.g., [121, 125, 72, 54, 35, 18]). Despite this extensive literature from an end-user perspective, few studies have considered smart home product teams. This is a notable gap, as the perceptions, perspectives, and practices of those designing and developing smart home devices may play a crucial role in multi-user smart home privacy. Although some work has been conducted with smart home designers [25, 26, 11, 12], there remains a gap in understanding how smart home product teams—comprising software/hardware engineers, product managers, legal/regulatory compliance advisors, and marketing staff—address privacy-related challenges.
Furthermore, the current body of research on smart home privacy is geographically focused on WEIRD (Western, Educated, Industrialized, Rich, Democratic) countries [78, 25, 26, 89]. Smart home privacy in non-WEIRD countries has received limited scholarly attention [45]. As a non-WEIRD nation, China offers a distinct set of social, cultural, and legal characteristics that can surface novel privacy challenges and design considerations in the realm of smart home technologies [47], including: i) the country’s accelerated digital transformation [31]; ii) the high rate of smart home device adoption, with over 78 million Chinese households using such devices in 2023 [95]; iii) the large number of smart home product teams in China [92]; iv) the introduction of new privacy-focused laws and regulations within the country (§2.4); and v) China’s tech landscape—driven by firms like Xiaomi instead of Western giants—which has given rise to alternative data governance practices and innovation pathways [128].
Our study is the first to examine the privacy views and practices of Chinese smart home product team members, focusing on those involved in designing, developing, or evaluating smart home devices. These teams typically include software/hardware engineers, UX designers, product/project managers, legal/privacy experts, and marketing or operations specialists. Our research questions (RQs) are as follows:
@@
RQ1. Participants highlighted the importance of legal/ regulatory compliance in the design and development of smart home devices, often participating in routine reviews and showing a strong ethical commitment to user privacy. However, many—particularly UX designers—lacked adequate knowledge of privacy laws/data practices and formal training. They also noted that China’s unique sociocultural context (e.g., collectivism and Confucian values) contributed to lower privacy concern among Chinese smart home device users and limited legal literacy among product teams, shaping the design of privacy features.
@@
RQ2. Participants encountered several challenges, including: (1) managing complex and dynamic relationships between government authorities, companies, and users – for example, situations where law enforcement agencies could access user data, bypassing internal privacy safeguards within companies; (2) limited attention to the privacy needs of at-risk groups (e.g., older adults, domestic workers), stemming from difficulties in conducting empirical research with these populations and ethical limitations; and (3) the need to balance users’ privacy with UX, such as ease of use and convenience, particularly within the shared environment of multi-user smart homes.
@@
RQ3. To address the challenges identified above, participants used various privacy practices and risk mitigation strategies. These included technical measures such as restricting data access and implementing decentralized data processing. They also advocated for social initiatives to improve user privacy awareness (e.g., community-level outreach activities) and called for clearer legal guidelines for smart home product design/development. Although participants acknowledged recent improvements in local privacy legislation, they highlighted a gap between legal frameworks and actual development practices, often caused by the fast-paced evolution of smart home technologies in China.
@@
Contributions. We conducted the first exploratory interview study focusing on Chinese smart home product team members to investigate their perspectives and practices related to privacy. Unlike prior studies, our participant sample includes smart home product team members in various roles mainly from large companies, providing a more comprehensive overview of their privacy perspectives and practices (see Table I). Additionally, this study is the first to explore how these product team members navigate and implement privacy laws and ensure legal compliance during smart home product design and development. Our findings provide valuable and rich empirical insights into privacy practices in the Chinese smart home industry, with implications for product teams, researchers, and policy makers, especially in light of China’s distinct sociocultural environment. Based on our empirical evidence, we extend the discussion on Privacy by Design by emphasizing the need for responsible and transparent guidelines adapted to China’s unique sociocultural context—such as Confucian values, collectivist norms, and China’s culturally specific understanding of “privacy” (§5.3). Further, our findings highlight disparities and power dynamics between large companies—which typically have greater resources, dedicated compliance teams, and structured internal legal support—and small- and medium-sized enterprises (SMEs), which often face significant challenges in privacy compliance due to resource constraints and limited legal expertise. Based on our empirical evidence, we strongly recommend clearer legal frameworks to guide privacy practices, specifically designed to support compliance efforts in SMEs. Finally, we call for enhanced awareness of privacy issues among both the general public and product team members through educational initiatives in communities and schools, as well as targeted training for legal/regulatory compliance and data management (§5.4).
@@ 2.1 Multi-User Smart Home Privacy
Previous studies have explored the privacy, security, and safety needs, preferences, and concerns of smart home device owners and primary users [1, 2, 121, 107, 125, 98, 63, 54, 72], including Chinese users [65, 53, 46], emphasizing the importance of data ownership by users, data practice transparency, and informed user consent. The studies have also highlighted how social norms, including privacy norms, along with factors such as power dynamics, social structures, demographics, and technical expertise, influence views on smart home device adoption and use [2, 15, 98, 93]. Other papers have explored the privacy implications of smart home devices for different bystander groups who share a living space, including visiting, live-in, and uninvolved bystanders [86], highlighting privacy and security concerns of at-risk groups such as domestic workers [17, 18, 93, 35, 3, 47] and household members living in hostile environments [14, 64]. Additionally, some studies engaged device owners/users and bystanders in the co-design and evaluation of prototypes for select design interventions proposed in the academic literature [72, 43, 117, 107, 73, 63, 7, 81, 122].
Building on the findings of the aforementioned studies involving primarily device owners and bystanders, researchers proposed technical solutions to improve multi-user smart home privacy, including minimizing data collection for essential functions [72]; selectively limiting data collection in sensitive situations [30, 43]; designing privacy nudges and notifications [39, 63, 125]; implementing temporary data storage mechanisms [43, 30]; using smart home dashboards for data visualization [72, 117, 107]; creating separate user profiles (e.g., guest profiles with limited data access and control) [14, 39, 54]; employing tangible practices (e.g., camera lens caps) [7, 6]; and implementing communication and negotiation mechanisms to state privacy preferences [43, 10, 131]. Other studies proposed social interventions to address privacy concerns in multi-user smart homes, e.g., design guidelines aimed at supporting smart home product teams [17, 25, 10].
@@ 2.2 Privacy in Technology Design/Development
Previous research on privacy in digital technology design and development focused primarily on the challenges encountered by software developers of mobile apps in eliciting and incorporating privacy-related requirements. These challenges include privacy- and security-related issues in development [66, 103, 57, 88, 4], the failure to comply with privacy laws and regulations [13, 102], and guidance documentation that inadvertently leads developers to privacy-compromising decisions [104, 59]. In addition, privacy concerns have been identified in APIs and tools used by programmers [29, 42]. Research also showed that developers often deprioritize privacy due to stakeholder pressures [36, 59], placing greater emphasis on revenue generation and functionality than on privacy protection  [105, 58, 99]. (Refer to [78] for a review of the factors influencing software developers’ security and privacy practices.) Unlike prior studies, our paper investigates the privacy perceptions and practices of Chinese smart home product teams, thereby extending the focus from software developers to a broader range of product team members in diverse roles and organizations, focusing on smart home devices.
@@
In smart homes, Chalhoub et al. examined how product teams working in UK companies integrate UX principles into the security and privacy design of smart home cameras, proposing solutions such as customizable privacy settings and transparent user consent mechanisms [25, 26]. These studies were conducted in a WEIRD context, focusing exclusively on a single type of smart home device—smart cameras—and on end-user experience. They did not study how product teams perceived the privacy concerns of other stakeholders in smart homes, including bystanders, and did not take into account the power dynamics that exist among these groups. Zhang and Xie examined the main activities of and challenges faced by IoT developers—mainly hardware engineers, embedded software developers, and app developers—in Chinese (non-WEIRD) start-ups [130]. Although the study did not focus on security and privacy in the smart home context, and recruited participants only from startups, it found that participants frequently overlooked security issues due to limited resources and skills.
@@
Finally, recent work [9, 10] showed how socioeconomic factors, social norms, and religious beliefs influence power dynamics in Jordanian (non-WEIRD) smart homes. However, these studies did not explore the privacy perceptions and practices of local smart home product teams, focusing instead on end-users and a bystander group: domestic workers. While follow-up work [11] focused on smart home designers and their lack of privacy-support tools and guidelines, our study offers novel insights into Chinese product teams’ perspectives and practices, examining privacy considerations across multiple roles. Additionally, our research highlights how cultural values (e.g., filial piety), priorities (e.g., prioritizing legal compliance over individual privacy), and hierarchical, collectivist dynamics influence privacy trade-offs within these teams.
@@ 2.3 Confucianism and Privacy in China
Confucianism, a foundational philosophy in Chinese culture, emphasizes social harmony, hierarchical relationships, and the importance of family and community [120]. Central to Confucian thought are concepts like filial piety (“xiao”), which underscores respect and duty toward one’s parents and elderly people, and collectivism, which prioritizes group needs over individual desires [49]. Some studies have highlighted that traditional Confucian thinking, which significantly affects the understanding of privacy within the Chinese family unit, is often associated with the indigenous Chinese concept of “yin si” with notions of unspeakable immorality, indecency, or shame, providing minimal support for the development of legal norms that favor individual privacy protection [76, 114, 124]. Consequently, privacy norms in China have developed at a slower pace compared to Western societies, where privacy is deeply rooted in concepts of liberal individualism [16] and central to liberal political theory [48].
2.4 Legal and Regulatory Frameworks in China
@@
In China, companies’ privacy practices are primarily governed by the Personal Information Protection Law (PIPL) [^1] Personal Information Protection Law (PIPL), Cybersecurity Law (CSL) [^2] Cybersecurity Law (CSL) of the People’s Republic of China, Data Security Law (DSL) [^3] Data Security Law (DSL) of the People’s Republic of China, and Provisions on Online Protection of Children’s Personal Information (POPCPI) [^4] Provisions on Online Protection of Children’s Personal Information. These Chinese legal frameworks establish data privacy protections and define the scope of state oversight over data processors and companies [56, 112]. While GDPR, CCPA, and PIPL share common goals of protecting personal data and individual rights, they differ in their enforcement mechanisms, the extent of government access, and overall emphasis. Although PIPL is influenced by GDPR and incorporates similar individual rights, it places greater emphasis on national security and allows broader government access to data [23, 38, 111]. CSL and DSL contain provisions related to personal information protection but they both prioritize national security over individual privacy rights [23]. In contrast, PIPL focuses on individuals’ privacy within the Chinese context (e.g., Chapter 3). However, certain terms in PIPL—particularly those related to separate consent and the criteria for lawful anonymization—are open to interpretation, creating uncertainty in their application [123, 111]. PIPL supports individual data protection, but allows government access under certain circumstances (e.g., Chapter 2). Furthermore, enforcing data protection laws in China is complex due to the involvement of multiple administrative authorities, overlapping responsibilities, inconsistent enforcement practices, and internal contradictions within the legal framework [32, 129].
@@
Age Gender Education Experience Location Role Firm size Type Designed/Developed device(s)
@@
P1-EN 25-34 Man Master’s 6-10 yrs Shanghai Hardware engineer 1000+ MNC Smart health devices/monitors
@@
P2-PM 25-34 Man Master’s 1-5 yrs Shanghai Product manager 1000+ MNC Smart speakers, Smart home cameras
@@
P3-PM 25-34 Man Bachelor’s 6-10 yrs Shanghai Product manager 11-50 PE Baby cameras/monitors, Smart lighting
@@
P4-PM 25-34 Man Bachelor’s 1-5 yrs Guangdong Product manager 1000+ PE Baby cameras/monitors, Smart home cameras, Door locks/doorbells
@@
P5-UX 25-34 Man Bachelor’s 10+ yrs Shanghai UX design manager 1000+ PE Smart speakers, Smart home cameras, Door locks/doorbells, Smart thermostats, Smart TVs, Smart lighting, Smart plugs
P6-UX 25-34 Woman Master’s 6-10 yrs Sichuan UX designer 1000+ SOE Smart speakers, Smart TVs
@@
P7-PM 25-34 Man Master’s 1-5 yrs Guangdong Product manager 501-1000 PE Smart speakers, Door locks/doorbells
@@
P8-UX 25-34 Woman Master’s 6-10 yrs Shanghai UX designer 1000+ MNC Smart health devices/monitors, Smart thermostats
@@
P9-UX 25-34 Man Master’s 1-5 yrs Chongqing UX designer 1000+ SOE Smart speakers, Smart home cameras, Smart thermostats
@@
P10-PM 35-44 Man Bachelor’s 10+ yrs Jiangsu Product manager/director 1000+ PE Smart speakers, Smart home cameras, Door locks/doorbells, Smart health devices/monitors, Smart thermostats, Baby cameras/monitors, Smart lighting, Smart plugs, Smoke detectors/alarms
P11-PM 25-34 Man Master’s 6-10 yrs Zhejiang Project manager 1000+ SOE Smart home cameras
@@
P12-PM 25-34 Woman Bachelor’s 1-5 yrs Beijing Product manager 1000+ PE Smart home cameras, Door locks/doorbells, Baby cameras/monitors
@@
P13-UX 25-34 Man Master’s 6-10 yrs Zhejiang UX designer 1000+ SOE Smart home cameras, Door locks/doorbells, Smoke detectors/alarms
@@
P14-EN 25-34 Man Bachelor’s 6-10 yrs Guangdong Sr. software engineer 51-100 PE Smart speakers, Smart home cameras, Door locks/doorbells, Smart health devices/monitors, Smart Thermostats, Smart lighting, Smart plugs, Smoke detectors/alarms, Kitchen appliances
@@
P15-EN 35-44 Man Master’s 10+ yrs Shannxi Software director 1000+ SOE Smart health devices/monitors, Smart TVs, Kitchen appliances
@@
P16-PM 25-34 Woman Master’s 1-5 yrs Guangdong Product manager 0-10 PE Smart speakers, Smart home robots
@@
P17-EN 25-34 Man Doctoral 1-5 yrs Beijing Algorithm engineer 501-1000 PE Smart watches, Smart home cameras
@@
P18-LE 25-34 Man Master’s 1-5 yrs Beijing Safety compliance engineer 1000+ PE Smart TVs, Smart Thermostats, Smart lighting, Smart speakers, Smart plugs, Smart home cameras, Door locks/doorbells, Smart health devices/monitors, Smoke detectors/alarms
@@
P19-LE 25-34 Woman Master’s 1-5 yrs Shanghai Compliance officer 1000+ PE Smart home robots
@@
P20-LE 25-34 Man Bachelor’s 6-10 yrs Guangdong Product compliance engineer 101-500 PE Smart plugs, Smart home cameras, Door locks/doorbells, Kitchen appliances
P21-EN 25-34 Man Master’s 6-10 yrs Guangdong Software engineer 1000+ SOE Smart home cameras
@@
P22-EN 25-34 Woman Bachelor’s 1-5 yrs Jiangsu Software engineer 1000+ PE Kitchen appliances, Smart home robots
@@
P23-LE 25-34 Woman Master’s 1-5 yrs Zhejiang Compliance officer 1000+ SOE Smart speakers, Smart home cameras, Baby cameras/monitors, Kitchen appliances
@@
P24-EN 35-44 Woman Bachelor’s 10+ yrs Chongqing Software director 1000+ SOE Smart thermostats, Smart speakers, Smart home cameras
@@
P25-MS 25-34 Woman Master’s 6-10 yrs Shanghai Marketing executive 101-500 MNC Smart plugs, Kitchen appliances
@@
P26-MS 25-34 Woman Master’s 6-10 yrs Shanghai Digital marketing specialist 51-100 PE Smart home cameras